OpenStudy (anonymous):

Worker out corner has 5 more than 3times as many execise bicycles as the gym.Together they have 21 bicycles. slove the equation x+3x+5=21 to find the number of bicycles at work out corner.

OpenStudy (anonymous):

help

OpenStudy (anonymous):

hello can someone hel me

OpenStudy (anonymous):

4x+5=21

OpenStudy (anonymous):

4x+5-5=21-5

OpenStudy (anonymous):

4x=16

OpenStudy (anonymous):

\[\frac{ 4x }{ 4}=\frac{ 16 }{ 4 }\]

OpenStudy (anonymous):

x=?

OpenStudy (anonymous):

then workout corner has 5+3x

OpenStudy (anonymous):

5 more than 3 times the amount